A step-down transformer has a turn's ratio of 1:20
Answer #1

The tension in the secondary is

V_S=V_P\frac{N_S}{N_P}=20V_P=2400V

Thus

P=\frac{V_S^2}{R}=19200W

I'm not sure if I correctly understood the turns ratio: if I w as wronf andthe secondary has 1/20 of the primary turns then

V_S=V_P\frac{N_S}{N_P}=\frac{V_P}{20}=6V

and power is

P=\frac{V_S^2}{R}=0.12W